Clifford, Inc., has a target debt-equity ratio of 1.20. Its WACC is 8.7 percent, and the tax rate is 22 percent. a. If the company’s cost of equity is 13 percent, what is its pretax cost of debt? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) b. If instead you know that the aftertax cost of debt is 7.1 percent, what is the cost of equity?
Answer #1

Debt-equity ratio=debt/equity

Hence debt=1.2*equity

Let equity be $x

Debt=$1.2x

Total=$2.2x

WACC=Respective costs*Respective weight

a.

8.7=(x/2.2x*13)+(1.2x/2.2x*Cost of debt)

8.7=5.9090909+(1.2x/2.2x*Cost of debt)

Cost of debt=(8.7-5.90909091)*(2.2/1.2)

=5.1167%(Approx)

Hence pretax cost of debt=cost of debt/(1-tax rate)

=5.1167/(1-0.22)

=6.56%(Approx).

b.

8.7=(x/2.2x*Cost of equity)+(1.2x/2.2x*7.1)

8.7=(x/2.2x*Cost of equity)+3.87272727

Cost of equity=(8.7-3.87272727)*2.2

=10.62%